Find out aggregate income and aggregate consumption when investment increases by 100 crore in 2008.
Year 2006
Aggregate income 10000
Aggregate consumption 9000

Year 2007
Aggregate income 11000
Aggregate consumption 9900

Year 2008
Aggregate income ?
Aggregate consumption ?
1
Expert's answer
2020-11-23T08:33:28-0500

An increase in investment will cause a rise in output thus increasing income. Therefore, in 2008, income rises by the same rate as investment.

Aggregate Income 2008=aggregate Income 2007+100

=11000+100=11100

An increase in investment also directly affects productivity causing an increase in demand and an equal rise in consumption.

Aggregate consumption in 2008=Aggregate consumption in 2007+ 100

=9900+100=10000